Project

General

Profile

NDOS Commissioning Meeting 2011/02/28

Attendance: Jon, Tricia, Andrew, Mark Messier, Mark Mathis, Enrique, Minerba, Gary, Alena, Tom, Leon, Rick K., Greg, Ralf, Kurt, Denis, Mat, Art

Agenda:

  1. Channel Recovery Plans
  2. DAQ Update & Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

Channel Recovery:

  • Minerba: Retuning code to optimize use of threshold
  • Jon: need a database table to track movement of hardware (eg, APDs, FEBs) in and out of detector
  • Leon:
    • planning on testing cutting of ribbon cable today to see
    • hardware accounting:
      • we have 115 APDs + mounting hardware of current vintage
      • about 15 were part of earlier prototype run which had undersized mounting holes. These could probably make use of these by using turn-pins.
      • 80 for next di-block, 33 for muon catcher. Doesn't leave much wiggle room.
      • could possibly instrument less of muon catcher.
    • Jon: time scale for getting all of equipment here? Leon: will need ~30 arrays to commissioning APD test stand
    • DCM status: one spare is nearly ready, there is a short on the board that is being tracked down.
    • 3rd di-block is not filled yet, won't be filled until
  • Mat: seven of the oily APDs have been tested and seem to still be good after several tests. Would like to install these on the detector. 5 empty spots on top of diblock 2. Waiting for heat-sinks.
  • Mark Messier: status of noise-killing boards? Leon: should find out more tomorrow.
  • Jon: will follow up with Rick, Leon to come up with a plan for the rest of the month.

DAQ Software:

  • Kurt: latest version of software will automatically populate dB with DAQ/DCM configuration.
  • Leon: it would be good to get this information into the DSO analysis scripts.
  • Andrew: last Thursday there was an incident where the MTU on two of the event-builder buffer nodes were modified without knowledge or consent
  • Bus arbitration bit setting could improve performance of DCMs by factor of ~5.
  • Sue would like some time to expand size of Ganglia database. Could lose up to 30 minutes of data. Will coordinate with the shifter to do this during the next DSO scan (2 pm).

DSO data:

  • Leon: close to automating process of determining pixel masks and thresholds
  • Denis: will work on getting code to do this into the PedestalRunner.
  • Jon: any idea why cycling power of DCMs helps speed up DSO scans? Andrew: not really. DCMs FPGA or kernel module not being correctly reset?
  • Jon: any progress on parallelizing DSO scan? Andrew: hasn't had time to dig further.

DCS (Andrew): some hardware work was done last week that left a mess of cables out at the detector. This poses a danger, and is unacceptable.